VEGAS Team Meeting

Attendees: Amanda, Anish, Dan, Jack, Paul, Joe, Marty, Richard, Mark, John

Agenda:

  1. DONE Status, plans and priorities for ADC calibration
  2. DONE Status of INTEGRAT testing
  3. DONE Status of Mode 10-19 (ghost frequencies) issue
  4. Gather, and prioritize maintenance time requests. Maintenance days will be Thursday (all day) and Friday (till 2p)
  5. Status of shared-risk astronomical observations: https://safe.nrao.edu/wiki/bin/view/CICADA/CICADAGreenBankSpectrometerAstronomicalObservations
  6. Status of other “current issues”: https://safe.nrao.edu/wiki/bin/view/CICADA/CICADAGreenBankSpectrometerCurrentIssues

Meeting notes:

  1. Status, plans and priorities for ADC calibration - Three open issues
    1. OGP calibration followed by clock rate change yields different calibration values - no ideas yet, Jack & Hon were involved in discussions but we should add Dave McMahon -
      • The phase needs to change as a function of clock frequency - does this indicate a ongoing need to run the OGP calibration?
      • This would be loaded from a table, not recalculated.
    2. Certain bof/rate combinations fail to deliver a usable calibration parameter
      • This should not happen - Jack is not seeing these issues in his local ADCs. He will see if he can recreate the problem there.
      • Paul will send the details of combinations that cause the problem and respond to most recent bof compiling question from Jack
      • Seems to happen to one "zdoc" of the two - not sure if it is consistently one or two. (Jack has only seen it on 'one')
    3. OGP still fails one in for four times but this is okay for a "maintenance" function
    4. Storing MMCM values in FITS files. Does OGP/INL need to be written as well? - no, Paul will email gbsapp
  • This has been a really long progress to resolve problem, but approaching the final solution
  1. Status of INTEGRAT testing
    • Needs additional testing before closure
  2. Status of Mode 10-19 (ghost frequencies) issue
    • Not much change because Glenn unavailable for a couple of weeks
    • L8/drop 7 subbands workaround works for all but a few higher channel modes - can be accommodated with observation-specific scripts to use the workaround
      • Randy can look at firmware code and see if that helps and then determine if we need software assistance.
  3. Weekly testing schedule
    • Tuesday
      • LO1 Installation
    • Wednesday
      • 8:00 - 11:00 Test ADC calibrations (Amanda)
    • Thursday
      • Am open
      • 1:00 - 3:00 Collect data for INTEGRAT fix (Amanda)
      • 3:00 - 4:00 User config testing (Amanda)
  4. Status of shared-risk astronomical observations: https://safe.nrao.edu/wiki/bin/view/CICADA/CICADAGreenBankSpectrometerAstronomicalObservations
  5. Status of other “current issues”: https://safe.nrao.edu/wiki/bin/view/CICADA/CICADAGreenBankSpectrometerCurrentIssues

Other

  1. Please make sure to properly code your timesheets for the current time period and going forward.
    • Software engineers should be using a GBT ops software CM&E account - Account 893252
    • Electronic engineers should be using a GBT ops electronics CM&E account - Account 432510
    • Commissioning activities should be getting charged to hardware support and testing - Account 432510
    • Scientists assisting with shared-risk commissioning should be charging an observer support account
      • 432540 for the most part. Accounts 142020 (GB-based staff) and 142010 (CV-based staff) for when you are helping with overarching observing support (not problem solving, VEGAS specific)

-- MartyBloss - 2014-06-02
Topic revision: r1 - 2014-06-02, MartyBloss
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback